OpenStudy (anonymous):

A soap bubble had rainbow-colors and broke after forming a colorless part. What is the name of “colorless part”? Calculate its thickness as “a water film (n = 1.333) in air (n = 1)”. [*Visible range is 400 – 800 nm.]

Awolflover1 (awolflover1):

Options?

OpenStudy (anonymous):

A soap bubble had rainbow-colors and broke after forming a colorless part. What is the name of “colorless part”? Calculate its thickness as “a water film (n = 1.333) in air (n = 1)”. [*Visible range is 400 – 800 nm.] no option is given

OpenStudy (farcher):

This question is about thin film interference. Like from outside the bubble is reflected from the outside surface and the inside surface of the bubble. These two reflected beams overlap and hence interfere with one another. The analysis of what happens is complicated by two facts. The light of wavelength lambda in the air has a wavelength of lambda divided by the refractive index of water in the bubble (which is mainly water). There is a 180 degree (pi) phase shift when light is reflected at the outside surface of the bubble. The condition for a maximum for given wavelength lambda is \[2 t = (m-\frac 1 2) \frac \lambda n\] where t is the thickness of the film, n is the refractive index of water and m is a positive integer. You will note that as the thickness of the film gets smaller the value of m to get a maximum must get smaller. The smallest possible value of m is 1 and so the minimum thickness of a film so that a reflection occurs is \[t_\textrm{\min} = \frac {\lambda}{4n}\] As the thickness of the film decreases first red light (with the longest wavelength) is no longer reflected, the orange etc until the film is so thin that even the violet light is no longer reflected. So there is no reflected light from the film as the film gets thinner and thinner beyond this thickness.